# An All-Optical General-Purpose CPU and Optical Computer Architecture

### v1

@@ I Introduction
There is a very sensible push in industry and research, that the focus should not be on the computation itself, but rather on the communication between compute elements[8]. Currently, the energy consumption of electronic communication happening on and off chip is over $80\%$ of the total power consumption of any electronic processor. By replacing these electronic wires with optical interconnects, one can expect an immediate efficiency gain of up to a factor of $6$, due to waveguides being near lossless and highly efficient electro-optical modulation schemes. The industry expects the majority of efficiency gains in the near future coming from this electro-optical hybrid computing approach, with both large companies, like Intel and Nvidia, as well as startups, like Lightmatter, Ayar Labs, Celestial AI and Black Semiconductor, actively developing schemes for optical interposers to connect electronic computing elements and memory. This development is highly beneficial not just for the field of computing, but also photonics in general.
@@
With our research, however, we are focused on the phase thereafter. Once optical interconnects and interposers have been fully established and are the main mode of inter-chip and intra-chip communication, solving the 6×6$\times$6 × inefficiencies, the bottleneck will shift once again. By essentially eliminating the inefficiencies from communication itself, electronic computation and memory will become the main source of power consumption in excess of $90\%$, with electro-optical signal conversion accounting for the majority of the remaining percentage. Our focus, thus, is on reducing the energy consumption of compute and memory through an all-optical approach, to unlock another 10×10$\times$10 × efficiency increase in the long term.
@@
The paper introduces the required computer architecture concepts as needed and does not assume prior knowledge in the field.
@@ I-A Digital Optical Computing Past and Present
TABLE I: Miller’s criteria for practical optical logic reproducedfrom [15] (shortened).
Criteria
Description
Cascadeability
The output of one stage must be in the correct form to drive the input of the next stage.
essential
Fan-out
The output of one stage must be sufficient to drive the inputs of at least two subsequent stages (fan-out or signal gain of at least two).
Logic-level restoration
The quality of the logic signal is restored so that degradations in signal quality do not propagate through the system; that is, the signal is ‘cleaned up’ at each stage.
Input/output isolation
We do not want signals reflected back into the output to behave as if they were input signals, as this makes system design very difficult.
Absence of critical biasing
We do not want to have to set the operating point of each device to a high level of precision.
optional
Logic level independent of loss
The logic level represented in a signal should not depend on transmission loss, as this loss can vary for different paths in a system.
@@
The almost canonical way of performing all-optical switching and logic is to use semiconductor optical amplifiers (SOA) and exploit their cross-gain modulation (XGM) or cross-phase modulation (XPM) capabilities[21]. With very reliable devices having been shown over the past 20 years[22], SOAs have proven useful for various types of all-optical operation, including decoder logic[23, 24] and signal regeneration[25, 26, 27]. The recovery time of the SOA limits its performance, but it has been shown that more than 320 Gbit/s[28] all-optical switching is possible, with some implementations enabling even the Tbit/s domain[29]. While SOAs are mostly based on III-V semiconductor technologies, such as Indium Phosphide (InP), efforts are being made to integrate them with silicon photonics, including with a focus on optical computing[30]. But there are downsides to using SOAs, such as high power consumption and heat output, as well as integration cost. To address these issues for use in optical computing, alternative materials are explored, most notable 2D-materials[31]. The strong optical nonlinearities in 2D materials lend themselves well for all-optical switching[32] and these all-optical switches can be integrated directly with the common Silicon-on-Insulator (SOI)[33, 34] and Silicon Nitride (SiN)[35, 36] photonics platforms. The nonlinearities explored in graphene, such as saturable absorption, allow for extremely fast switching[37, 35], with recovery times well below $<1ps$, offering a path forward from SOAs.
The discussion on memory is twofold. On the one hand, we need an element that is able to store information to be read (and ideally written to) optically, on the other hand we require a decoding scheme to address the individual bits. For a comprehensive review on optical memory technologies, we refer to Lian etal. [38] and Alexoudi etal.[39]. The topic of addressing memories is less well explored. While the previous all-optical switching devices do lend themselves to be used this way, it requires considerable power for large address spaces[24]. Instead, a more elegant approach is to use passive devices as optical decoders for addressing memory[40, 41].
@@ II-A Efficient General-purpose without billions of transistors
General-purpose computing merely reflects the fact, that the device is itself able to fully simulate a Turing machine (TM)[46], given an infinite amount of memory. For obvious reasons, here we will refer by TM to a device that is practically Turing-complete (TC) but does not have an infinite amount of memory. In turn, the TM itself is the simplest computer able to run Doom. While such a TM would be easily implemented using a handful of logic elements as an all-optical computer, it is highly impractical and not easily compatible with classical software. Instead, we investigate the simplest TC RISC architecture that we can implement, namely SUBLEQ.
@@ II-F Architecture beyond von Neumann
The core of the XPU is the purely digital feed-forward logic processing unit (LPU), which contain no memory of their own and implement the typical features found in a general-purpose processor: instruction decoder, arithmetic logic unit (ALU), instruction pointer logic and memory controllers. All the typical registers one would find in a CPU are decoupled and external in our LPU. Multiple of these digital LPUs can populate the XPU. Each of these 16-bit units implement a reduced instruction set based on a load/store architecture and simple integer arithmetic, with one additional specialized instruction called reconfigurable functional unit call (RFU). In its simplest form, the SUBLEQ processor explored here can be used as an LPU, if extended by an RFU instruction and the ability to access all the different memory types.
@@ III-A All-Optical Logic
We presented a variety of all-optical decoding schemes in Section I-A. To test the feasibility of our approach, we implement a hybrid decoder using a mixture of the discussed XGM approach using SOAs in an InP platform and passive SiN elements, which both individually have been shown to be compact and fulfill a majority of the criteria laid out by Miller in similar applications. The resulting one-hot [^5] Alternative to binary encoding, where one bit is set to $1$ and all others to 00. The position of the bit defines the value. For example 101 (5) in binary is equivalent to 00100000 in one-hot. decoded digital signal allows us to perform arbitrary logic functions using a crossbar as shown in Figure 6. Usually, these crossbar configurations are used for analog processing, however by utilizing it in binary, the structure simplifies considerably. Essentially, the crossbar becomes a 1-to-1 reflection of the binary truth table, making it sparse depending on the logic implemented. This sparsity allows for much more compact designs, as the logical 0 values are not implemented. For the automated design pipeline (Figure 8), this is an important consideration, as in an all-optical setting, the canonical NAND gate is the most inefficient and to be avoided during logic synthesis.
@@
Figure 6: Working principle of the sparse crossbar in conjunction with an all-optical 3x8 decoder for digital logic. Here demonstrated 3 use-cases, as a full-adder with 2 outputs, a 3-input AND and a 3-input NAND gate. The orange and purple beams highlight the path of the individual look-ups.
@@
A further advantage of this structure of all-optical logic is, that multiple results can be generated in parallel, without the need for additional decoders or dedicated logic gates. This, for example, allows us to generate the output of a full-adder, which would require a multitude of logic gates, using a single decoder and crossbar, while still enabling us to output additional logic functions using that same input. Furthermore, using tuning elements or phase change materials, a logic function can be reconfigured if needed, much like an FPGA. Essentially each Decoder with a crossbar acts as a look-up table (LUT) or read-only memory (ROM). Another optimization to consider is the use of each crossbar using multiple wavelengths simultaneously (as hinted at by the purple and orange in Figure 6). Operating using multiple wavelengths would require a compatible decoder scheme. Since we are computing in binary and in a one-hot encoding, there is no need for exact control of interference pattern and merely “light or no light“, simplifying the design.
@@ III-C The CPU
Figure 9: Implementation of the first all-optical CPU. (a) Shows the steps the SUBLEQ-based processor needs to perform, where green highlights the logic operations, orange the memory access and blue the registers. Here, $M$ denotes the width of instruction ROM address space, $N$ the width of the memory address space and the *** indicates, that these memories are not needed to be uniform and can be of any type (RWORM, volatile, etc.). (b) Shows the photonic implementation of (a) using the same color scheme. Highlighted here also the physical PICs, left using a SiN platform for the fixed crossbar and heaters for fine tuning and right the PCM-enhanced SiN for the memory, both of which are connected to a separate non-linear element for decoding. To enable flexibility and debugging, the devices are interconnected using discrete fiber optics in this demonstration.
@@ III-D Results
All tests were conducted on bare PIC dies to allow quick debugging. To allow for testing of multiple devices per PIC, the PCM-based RWORM chips use grating couplers and are expected to have higher losses in this iteration, than one optimized for insertion loss. The configuration of the test setup can be seen in Figure 9.
@@
Considering that a typical CMOS full-adder comprises of $28T$ (transistors), the depicted crossbar configuration could be said to be equivalent to $\approx 40T_{e}$ in electronics. Our goal here is not to highlight the efficiency of all-optical logic implementations, but rather to point out the contrary: there is no value in comparing the abundant complexity of electronic circuit design with all-optical computing, as they follow completely different paradigms. A 1-to-1 comparison of integration density is not useful and undermines almost all of the advantages optical computing offers. For example, the $\approx 40T_{e}$ estimate does not include additional logic outputs in the crossbar, use of wavelength multiplexed mode or the ability to multithread reusing the same circuitry. These strategies are hard to quantify in transistor equivalent, as they do not scale linearly if realized in an electronic circuit. That being said, for our case, each logic unit was designed for 8 usable logic outputs, 1 wavelength and 40 parallel threads, and we will use a conservative $\approx 1500T_{e}$ equivalent for our discussion here.
@@ IV Conclusion and Outlook
We showed how the physical CPU presented here can be expanded on to be used as the logic processor (LPU) for the XPU computing architecture and gave a roadmap to realize a full XPU. Our efforts are now aimed towards the next iteration of the decoder and realizing it using 2D materials for a massively more energy efficient and faster design compared to SOAs. We are also focused on expanding the decoding and logic scheme to much wider ranges. This all will allow us to reach ultra-high bandwidth computing, ultra-low latency memory running at minimal power in the near future.

### v2

@@ I Introduction
There is a very sensible push in industry and research, that the focus should not be on the computation itself, but rather on the communication between compute elements[8]. Currently, the energy consumption of electronic communication happening on and off chip can reach over $80\%$ of the total power consumption of an electronic processor[9, 10, 11]. By replacing these electronic wires with optical interconnects, one can expect an immediate efficiency gain of up to a factor of $6$, due to waveguides being near lossless and highly efficient electro-optical modulation schemes. As an example, communicating a bit across an electronic chip can cost up to 600 fJ, whereas using electro-optical modulation would enable far lower energies, down to the 1-10 fJ domain[12], with the potential of reaching the attojoules in the coming years. The industry expects the majority of efficiency gains in the near future coming from such electro-optical hybrid computing approach, with both large companies, like Intel and Nvidia, as well as startups, like Lightmatter, Ayar Labs, Celestial AI and Black Semiconductor, actively developing schemes for optical interposers to connect electronic computing elements and memory. This development is highly beneficial not just for the field of computing, but also photonics in general.
@@
With our research, however, we are focused on the phase thereafter. Once optical interconnects and interposers have been fully established and are the main mode of inter-chip and intra-chip communication, solving the 6×6$\times$6 × inefficiencies, the bottleneck will shift once again. By essentially eliminating the inefficiencies from communication itself, electronic computation and memory will become the main source of power consumption. Using [8] as an example again, reducing the interconnect power consumption from 600 fJ to 1-10 fJ per bit, would then result in the compute (100 fJ for a floating point operation in this example) to dominate the energy equation. Our focus should be on reducing the energy consumption of compute and memory either after or in parallel to reducing interconnect power consumption. Through such a combined, all-optical approach, the potential to unlock another 10×10$\times$10 × efficiency increase in the long term.
@@
The paper introduces the required computer architecture concepts as needed and does not assume prior knowledge in the field. Our main contribution will be a novel architecture using a blend of old and new concepts (such as our RWORM, delay line based registers and all-optical LUT3 in section III) for all-optical computing on a photonic platform, that enables feasible general-purpose operation.
@@ I-A Digital Optical Computing Past and Present
TABLE I: Miller’s criteria for practical optical logic reproduced
from [19] (shortened).
Criteria Description
Cascadeability The output of one stage must be in the correct form to drive the input of the next stage. essential
Fan-out The output of one stage must be sufficient to drive the inputs of at least two subsequent stages (fan-out or signal gain of at least two).
Logic-level restoration The quality of the logic signal is restored so that degradations in signal quality do not propagate through the system; that is, the signal is ‘cleaned up’ at each stage.
Input/output isolation We do not want signals reflected back into the output to behave as if they were input signals, as this makes system design very difficult.
Absence of critical biasing We do not want to have to set the operating point of each device to a high level of precision. optional
Logic level independent of loss The logic level represented in a signal should not depend on transmission loss, as this loss can vary for different paths in a system.
@@
The almost canonical way of performing all-optical switching and logic is to use semiconductor optical amplifiers (SOA) and exploit their cross-gain modulation (XGM) or cross-phase modulation (XPM) capabilities[25]. With very reliable devices having been shown over the past 20 years[26], SOAs have proven useful for various types of all-optical operation, including decoders logic[27, 28]. Amplified spontaneous emission (ASE) has, however, limited the cascadability of such SOA based all-optical switches. To combat these shortcomings, signal regeneration[29, 30, 31] schemes were introduced, however increasing the power required to operate such devices. Further, the recovery time of the SOA limits its performance, but it has been shown that more than 320 Gbit/s[32] all-optical switching is possible, with some implementations enabling even the Tbit/s domain[33]. While SOAs are mostly based on III-V semiconductor technologies, such as Indium Phosphide (InP), efforts are being made to integrate them with silicon photonics, including with a focus on optical computing[34]. Apart from ASE, there are further downsides to using SOAs, such as high power consumption and heat output, as well as fabrication cost associated with the low volume production of III-V compared to CMOS in the current landscape. To address these issues for use in optical computing, alternative materials are explored, most notable 2D-materials[35]. The strong optical nonlinearities in 2D materials lend themselves well for all-optical switching[36] and these all-optical switches can be integrated directly with the common Silicon-on-Insulator (SOI)[37, 38] and Silicon Nitride (SiN)[39, 40] photonics platforms. The nonlinearities explored in graphene, such as saturable absorption, allow for extremely fast switching[41, 39], with recovery times well below $<1ps$, offering a path forward from SOAs.
The discussion on memory is twofold. On the one hand, we need an element that is able to store information to be read (and ideally written to) optically, on the other hand we require a decoding scheme to address the individual bits. For a comprehensive review on optical memory technologies, we refer to Lian etal. [42] and Alexoudi etal.[43]. The topic of addressing memories is less well explored. While the previous all-optical switching devices do lend themselves to be used this way, it requires considerable power for large address spaces[28]. Instead, a more elegant approach is to use passive devices as optical decoders for addressing memory[44, 45] and considerable progress is being made using phase-change materials (PCM) for optical memory with novel addressing schemes[46, 47, 48]. For an overview of the use of PCMs in photonics now and in the future, we refer to the review by Prabhathan etal. [49].
@@ II-A Efficient General-purpose without billions of transistors
General-purpose computing merely reflects the fact, that the device is itself able to fully simulate a Turing machine (TM)[54], given an infinite amount of memory. For obvious reasons, here we will refer by TM to a device that is practically Turing-complete (TC) but does not have an infinite amount of memory. In turn, the TM itself is the simplest computer able to run Doom. While such a TM would be easily implemented using a handful of logic elements as an all-optical computer, it is highly impractical and not easily compatible with classical software. Instead, we investigate the simplest TC RISC architecture that we can implement, namely SUBLEQ. It should be noted, that SUBLEQ acts merely as a toy example for this demonstration and discussion, the final architecture will make use of a true RISC architecture.
@@ II-F Architecture beyond von Neumann
The core of the XPU is the purely digital feed-forward logic processing unit (LPU), which contain no memory of their own and implement the typical features found in a general-purpose processor: instruction decoder, arithmetic logic unit (ALU), instruction pointer logic and memory controllers. All the typical registers one would find in a CPU are decoupled and external in our LPU. Multiple of these digital LPUs can populate the XPU. Each of these 16-bit units implement a reduced instruction set based on a load/store architecture and simple integer arithmetic, with one additional specialized instruction called reconfigurable functional unit call (RFU). In its simplest form, the SUBLEQ processor used as a toy example here could be used as an LPU, if extended by an RFU instruction and the ability to access all the different memory types. But in actual implementations of this architecture, more potent RISC ISAs will be implemented for the LPU.
@@ III-A All-Optical Logic
We presented a variety of all-optical decoding schemes in Section I-A. To test the feasibility of our approach, we implement a hybrid decoder using a mixture of the discussed XGM approach using SOAs in a monolithic InP platform and passive SiN elements for interconnects and tuning elements, which both individually have been shown to be compact and fulfill a majority of the criteria laid out by Miller in similar applications. Our construction largely follows the canonical signal-probe MZI setup[101] in a monolithic InP platform[102, 103]. To reduce the contributions of ASE, saturable absorbers in the same platform are introduced in a 2R configuration[30], leading to a trade-off between total power consumed by the SOAs and the resulting extinction ratio (ER). While the 2R setup enables ER improvements of $>4db$ for input signals with an ER $<4db$, achieving an output extinction ratio throughout the system of $>5db$ is sufficient for our purposes.
@@
The resulting one-hot [^5] Alternative to binary encoding, where one bit is set to $1$ and all others to 00. The position of the bit defines the value. For example 101 (5) in binary is equivalent to 00100000 in one-hot. decoded digital signal allows us to perform arbitrary logic functions using a crossbar as shown in Figure 6. Usually, these crossbar configurations are used for analog processing, however by utilizing it in binary, the structure simplifies considerably. Essentially, the crossbar becomes a 1-to-1 reflection of the binary truth table, as follows,
$\begin{split}f_{\text{dec}}([\text{Bit}_{0},\text{Bit}_{1},\text{Bit}_{2}])&=\vec{x}_{\text{1-hot}}\\ f_{\text{xbar}}(f_{\text{dec}}([\text{Bit}_{0},\text{Bit}_{1},\text{Bit}_{2}]))&=\textbf{M}_{\text{xbar}}\cdot\vec{x}_{\text{1-hot}}\end{split}$ (1)
where, when we use the orange example of Figure 6, with the 3 input bits being 0, 0, 1, i.e. $f_{\text{dec}}(\textbf{{011}})=[0,0,0,1,0,0,0,0]$, we get
$\left[\begin{array}[]{cccccccc}0&1&1&0&1&0&0&1\\ 0&0&0&1&0&1&1&1\\ 0&0&0&0&0&0&0&1\\ 1&1&1&1&1&1&1&0\end{array}\right]\left[\begin{array}[]{c}0\\ 0\\ 0\\ 1\\ 0\\ 0\\ 0\\ 0\\ \end{array}\right]=\left[\begin{array}[]{c}0\\ 1\\ 0\\ 1\\ \end{array}\right]$ (2)
The multitude of 00 entries in the matrix allows for much more compact designs, as the logical 00 values do not need to be physically implemented (i.e., there is no crossbar at that intersection). For the automated design pipeline (Figure 8), this is an important consideration, as in an all-optical setting, the canonical NAND gate is the most inefficient and to be avoided during logic synthesis, whereas an AND gate or direct implementation of a full-adder is very efficient.
@@
Figure 6: Working principle of the sparse crossbar in conjunction with an all-optical 3x8 decoder for digital logic. Here demonstrated 3 use-cases, as a full-adder with 2 outputs, a 3-input AND and a 3-input NAND gate. The orange and purple beams highlight the path of the individual look-ups. Also included is a view on the 2 different possibilities of implementing the crossbar, either as non-reconfigurable (permanently set at fabrication) or as reconfigurable (using a naive PCM implementation).
@@
A further advantage of this structure of all-optical logic is, that multiple results can be generated in parallel, without the need for additional decoders or dedicated logic gates. This, for example, allows us to generate the output of a full-adder, which would require a multitude of logic gates, using a single decoder and crossbar, while still enabling us to output additional logic functions using that same input. Furthermore, by making the matrix $\textbf{M}_{\text{xbar}}$ in Equation 1 tunable, for example through the use of PCM, a logic function can be reconfigured after fabrication if needed, much like an FPGA. Since the crossbar is essentially digital (only allowing 00/off or $1$/on values) and is only changed occasionally, if ever (comparable to burning an EEPROM for an update), it acts robustly against small errors in fabrication or material variation. A compact and robust, yet highly inefficient design is shown in Figure 6. Essentially each Decoder $f_{\texttt{dec}}$ (performing 1-hot encoding) coupled to a crossbar $f_{\texttt{xbar}}$ (linearly transforming the 1-hot encoding to the desired output) acts as a look-up table (LUT) or read-only memory (ROM).
Another optimization to consider is the use of each crossbar using multiple wavelengths simultaneously (as hinted at by the purple and orange in Figure 6). Since we are computing in binary and in a one-hot encoding, there is no need for exact control of interference pattern and merely “light or no light“, simplifying the design. But operating using multiple wavelengths would require a compatible decoder $f_{\texttt{dec}}$ scheme, which is highly non-trivial.
@@ III-C The CPU
Figure 9: Implementation of the first all-optical CPU. (a) Shows the steps the SUBLEQ-based processor needs to perform, where green highlights the logic operations, orange the memory access and blue the registers. Here, $M$ denotes the width of instruction ROM address space, $N$ the width of the memory address space and the * indicates, that these memories are not needed to be uniform and can be of any type (RWORM, volatile, etc.). (b) Shows the photonic implementation of (a) using the same color scheme. Highlighted here also the physical PICs, left using a SiN platform for the fixed crossbar and heaters for fine tuning and right the PCM-enhanced SiN for the memory, both of which are connected to a separate non-linear element for decoding. To enable flexibility and debugging, the devices are interconnected using discrete fiber optics in this demonstration.
@@ III-D Results
All tests were conducted on bare PIC dies to allow quick debugging. To allow for testing of multiple devices per PIC, the PCM-based RWORM chips use grating couplers and are expected to have higher losses in this iteration, than one optimized for insertion loss (see also Section III in regards to design choices for the crossbar). The configuration of the test setup can be seen in Figure 9, where we note this is an abstract view and multiple decoders and crossbars were fitted per InP and SiN chip respectively.
Each PIC contains multiple copies of the used structures, coupled individually with grating couplers for characterization. The structures for the SiN PICs are: Waveguides, bends, PCM covered waveguides, grating couplers, variable directional couplers, edge couplers, terminators, crossings, heaters. The structures for the InP PICs are: Waveguides, bends, edge couplers, 1x2 MMIs, SOAs. Apart from the PCM covered waveguides, all components were designed according to the respective foundry’s PDK and characterization of the PICs has concluded that sufficiently many operate within the given parameters of the PDK [^6] Due to non-disclosure agreements with the foundries, exact results can not be published, but access to the PDKs can be obtained directly from the foundries.. As indicated in Section III-B, the results on the custom PCM-based components will be reported seperately, keeping the focus on the architecture here.
@@
Considering that a typical CMOS full-adder comprises of $28T$ (transistors), the depicted crossbar configuration could be said to be equivalent to $\approx 40T_{e}$ in electronics, factoring in the additional outputs. Our goal here is not to highlight the efficiency of all-optical logic implementations, but rather to point out the contrary: there is no value in comparing the abundant complexity of electronic circuit design with all-optical computing, as they follow completely different paradigms. A 1-to-1 comparison of integration density is not useful and undermines almost all of the advantages optical computing offers. For example, the $\approx 40T_{e}$ estimate does not include additional logic outputs in the crossbar, use of wavelength multiplexed mode or the ability to multithread reusing the same circuitry. These strategies are hard to quantify in transistor equivalent, as they do not scale linearly if realized in an electronic circuit. That being said, for our case, each logic unit was designed for 8 usable logic outputs, 1 wavelength and 40 parallel threads, and we will use a conservative $\approx 1500T_{e}$ equivalent for our discussion here.
@@ IV Conclusion and Outlook
We showed how the physical CPU presented here can be expanded on to be used as the logic processor (LPU) for the XPU computing architecture and gave a roadmap to realize a full XPU with a true RISC ISA instead of SUBLEQ. Our efforts are now aimed towards the next iteration of the decoder and realizing it using 2D materials for a massively more energy efficient and faster design compared to SOAs. We are also focused on expanding the decoding and logic scheme to much wider ranges. This all will allow us to reach ultra-high bandwidth computing, ultra-low latency memory running at minimal power in the near future.